OpenStudy (anonymous):

Solve 6 (sqrt x+1)-3=15 Answers 8 71 75 107

OpenStudy (anonymous):

first of all remove the -3 from the left hand side

OpenStudy (anonymous):

to get 6(sqrt x+1)=18

OpenStudy (anonymous):

then get rid of the 6 to get sqrt(x+1)=3

OpenStudy (anonymous):

the get rid of the root to get x+1 =9

OpenStudy (anonymous):

then get rid of the +1 to get x=8

OpenStudy (anonymous):

therefore the answer is 8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

haha no probs